Which of the following functions are solutions of the differential equation y”– 4y’+4y = e

x?

(a) e^x

(b) e^2x

(c) e^2x + e^x

(d) xe^2x + e^x

(e) e^2x + xe^x


1
Expert's answer
2021-09-16T08:07:13-0400

"\text{Using the method of undetermined co-efficients we find the particular solution to the}\\
\text{given differential equation.Let $y_p$ represent the particualar solution. }\\
\therefore y_p = Ae^t \text{ where A is a constant}\\
\implies y_p' = Ae^t, y_p'' = Ae^t\\
\text{Substituting into the given differential equation we have that}\\
Ae^t-4Ae^t+4Ae^t=e^t\\
\implies Ae^t = e^t\\
\therefore A = 1\\
\text{Hence, $e^t$ is a solution to the given differential equation.}"